The velocity of a particle moving along the x-axis varies according
1
Expert's answer
2019-09-25T09:02:53-0400

The velocity of a particle moving along the x-axis varies according to the expression

"v_x(t) = (40 – 5 t^2) \frac{m}{s}," where t is in seconds. Find the average acceleration in the time interval t=0 to t=2.0s.



"a_x=0-2(5)t=-10t"

The average acceleration in the time interval t=0 to t=2.0 s:


"a_{av}=0.5(a(0)+a(2))"

"a_{av}=0.5(-10(0)-10(2))=-10\frac{m}{s^2}"